OpenStudy (anonymous):

Your piggy bank has a total of 40 coins in it; some are dimes and some are nickels. If you have a total of $3.25, how many dimes do you have?

hero (hero):

x + y = 40 .05x + .10y = 3.25 y = 40 - x y = (3.25 - .05x)/.10 y = y .10(40 - x) = 3.25 - .05x 4 - .10x = 3.25 - .05x 4 - 3.25 = .05x .75 = .05x x = .75/.05 x = 15 y = 25

hero (hero):

@Luis_Rivera still doesn't work yet